OpenStudy (anonymous):

find the real valued solutions |x-3|=|5x|

jimthompson5910 (jim_thompson5910):

|x-3|=|5x| x-3=5x, or -(x-3)=5x, or x-3=-5x or -(x-3)=-5x Note: -(x-3)=5x is the same as x-3=-5x (after multiplying both sides by -1) also, -(x-3)=-5x is the same as x-3=5x (again, after multiplying both sides by -1) So we only have to solve 2 equations: x-3=5x and -(x-3)=5x to get the answers solve x-3=5x: x-3=5x -3=5x-x -3=4x -3/4 = x x = -3/4 solve -(x-3)=5x -(x-3)=5x -x+3=5x 3 = 5x+x 3 = 6x 3/6 = x 1/2 = x x= 1/2 So the answers are x = -3/4 or x = 1/2

OpenStudy (anonymous):

thank you :)